Assuming the length of drift tube in ToF MS is 2.3 m and accelerating voltage is 30,000 V, one type of ions can reach the ion collector in 11.3 µs. What is the molar mass of the ions, assuming the ions are singly charged cations?

Another instrumental analysis problem, please help!
Answer #1

The equation to find the mass(m) of a single ion is given by:

m=\frac{2qVt^2}{l^2}

where:

q - charge on ion = 1.6*10^(-19) C

V - voltage = 30,000 V

t - time taken = 11.3 µs = 11.3*10^(-6) s

l - length of drift tube = 2.3 m

Substituting the values, we get

m=\frac{2\times1.6\times 10^{-19}\times30000\times (11.3\times 10^{-6})^2 }{(2.3)^2}=2.3\times 10^{-25}\: kg

This is the mass of one ion. The molar mass is :

m=2.3\times 10^{-25}\: \times 6.022\times 10^{23}=14\times 10^{-2}\: kg=140\: g

Hence the molar mass of ion is 140 g/mol